### v2

# Generating Streamlining Constraints with Large Language Models

Abstract
Streamlining constraints (or streamliners, for short) narrow the search space, enhancing the speed and feasibility of solving complex constraint satisfaction problems. Traditionally, streamliners were crafted manually or generated through systematically combined atomic constraints with high-effort offline testing. Our approach utilizes the creativity of Large Language Models (LLMs) to propose effective streamliners for problems specified in the MiniZinc constraint programming language and integrates feedback to the LLM with quick empirical tests for validation. Evaluated across seven diverse constraint satisfaction problems, our method achieves substantial runtime reductions. We compare the results to obfuscated and disguised variants of the problem to see whether the results depend on LLM memorization. We also analyze whether longer off-line runs improve the quality of streamliners and whether the LLM can propose good combinations of streamliners.
1 Introduction
Streamliners are certain constraints added to a constraint model to reduce the search space, thereby improving the feasibility and speed of finding solutions to complex constraint satisfaction problems. By incorporating domain-specific knowledge, streamliners can guide the constraint solver, allowing it to bypass less promising areas of the search space. Gomes and Sellmann [?] introduced streamliners to speed up the constrained-based search for hard combinatorial design problems. Today, streamliners are a standard tool for speeding up constrained-based search. Streamliners are closely related to implied/redundant constraints, symmetry-breaking constraints, and dominance-breaking constraints; however, adding a streamliner may even cause the constraint model to become inconsistent.
Originally, streamliners were hand-crafted by researchers who used their theoretical insight to analyze the constrained model. However, progress has also been made on the automated generation of streamliners [?] by systematically trying the effect of some atomic constraints, such as imposing specific constraints on integer and function domains, like enforcing odd or even values, monotonicity, and properties like commutativity, as well as facilitating specific attributes in binary relations. These atomic restrictions are tested on thousands of problem instances, and those that show a good streamlining effect are systematically combined.
This paper proposes a different approach to automated streamliner generation that utilizes the creative power of Large Language Models (LLMs) to generate streamliners, providing a neurosymbolic prototype implementation StreamLLM, and rigorously tests it on seven different constraint satisfaction problems using at least 310 CPU days [^1] This is a conservative estimate; we spent about 24 CPU days for instance generation, 75 CPU days for grading the instances, 162 CPU days for evaluating the performance of the generated streamliners, 20 CPU days for the offline approach, and 28 days for the combinations approach.. Our approach leverages the capabilities of LLMs to infer potentially effective streamliners, similar to an experienced researcher’s intuitive grasp of a problem. By integrating LLMs into the streamliner generation process, we can potentially uncover unique and subtle patterns in the problem formulation and utilize them to speed up the solving process. To some extent, StreamLLM is closer related to a hand-crafted streamliner design by human experts rather than an automated bottom-up streamliner generation, the dominant strategy of the previous research [?].
Our system StreamLLM for streamliner generation combines queries to LLMs with quick empirical validation on small test instances solvable within seconds. The system generates effective constraints within minutes through engineered prompts and adaptive feedback based on streamliner performance. This realtime approach allows streamliner benefits to be realized even for hard problem instances that would typically require hours to solve—any significant speedup can outweigh the brief time spent on streamliner generation.
We evaluate StreamLLM extensively across seven diverse constraint problems, including standard benchmarks and a novel hypergraph coloring problem, using hundreds of instances specified in MiniZinc. The problems are formulated in the MiniZinc constraint programming language [?]. We test StreamLLM with two state-of-the-art LLMs (GPT-4o and Claude 3.5 Sonnet), different prompting variants, and different strategies for combining adaptive prompting with test runs on easy instances. Our results demonstrate remarkable effectiveness, with some streamliners reducing solving time by up to 99%.
To investigate whether these improvements stem from memorization or genuine problem understanding, we test StreamLLM on disguised and obfuscated variants of the problems. The system’s continued strong performance, particularly on problems with no known streamliners in the literature, suggests real analytical capabilities rather than mere pattern matching.
Beyond the realtime setting, we explore offline approaches with extended training periods and larger instance sets, achieving even better results in some cases. We also investigate StreamLLM’s ability to generate effective combinations of streamliners, finding that this often outperforms single constraints. Analysis of the generated streamliners reveals an interesting mix—some mirror expert-designed constraints while others take novel yet highly effective approaches.
Our experimental results suggest that LLM-generated streamliners can meaningfully reduce runtime across diverse constraint satisfaction problems with minimal problem-specific tuning, pointing to promising applications of AI assistance in constraint programming for computationally challenging problems.
Code and problem instances can be found on Zenodo [?].
2 Preliminaries
2.1 Constraint Programming
Constraint Programming (CP) is a methodology for solving combinatorial problems specified by means of declarative constraints. Please refer to Rossi, Van Beek, and Walsh [?] for a comprehensive discourse. Examples include scheduling, routing, planning, etc. These problems can be of two types – decision problems require a yes/no answer, while optimization problems require you to find a solution that minimizes or maximizes a given objective function. In this paper, we focus only on decision problems. MiniZinc [?] is a popular tool for solving CP problems. Problem specifications are written in the MiniZinc modeling language and are called models. MiniZinc compiles the model to a lower-level language and then solves it using one of many underlying solvers like Chuffed, Gecode, OR-Tools etc. Similarly, Conjure [??] is also a tool for higher-level constraint programming, which uses Essence as its modeling language. In this paper, we use MiniZinc as the CP solver of choice, while Conjure is used in the instance generation pipeline AutoIG. More specifically, the structure and format of the instances are specified in Essence and given to Conjure, which then generates instances following these constraints.
2.2 Streamliners
Streamlining constraints or streamliners, introduced by Gomes and Sellmann [?], are constraints that are added to the constraint programming models with the goal of narrowing the focus to a small but highly promising segment of the search space. Thus, they have the potential to significantly reduce the search effort for hard combinatorial problems. For example, Gomes and Sellmann [?] used constraints for “Latin squareness” as a streamliner for finding Diagonally Ordered Magic Squares. Streamliners have been successfully applied in a diverse range of settings222[????????????].
We note that streamlining constraints are not required to be sound. This means that adding the streamlining constraint may make a satisfiable instance of a constraint model unsatisfiable. As a consequence, an unsatisfiable instance in the streamlined model does not imply that the instance is unsatisfiable in the original model.
Other constraints that are similar to streamliners are implied (or redundant) constraints, symmetry-breaking constraints, and dominance-breaking constraints. Implied constraints do not change the set of feasible solutions333[?????]. Symmetry-breaking constraints eliminate certain solutions within each equivalence class while ensuring that at least one solution from each class remains444[??????]. Dominance-breaking constraints are applicable in the context of optimization problems (possibly formulated as a decision problem with the objective value explicitly stated in the model), as they disallow solutions that are known to be suboptimal. It might disallow optimal solutions as well, as long as at least one optimal solution remains555[????]. In contrast to streamliners, such constraints are guaranteed to be sound.
Finding a useful streamliner manually by a human expert is a time-consuming process. Therefore it is appealing to generate streamliners automatically, a direction explored successfully in previous work666[????]. A common point between all these previous approaches is that they treat the streamliner generation as a high-effort, offline task, taking up to 4 CPU days for a single problem. The streamliner is built up systematically from elementary constraints to reduce the domain of a decision variable and tested on a large set of automatically generated instances.
Similar to other recent studies [?] we will use the term streamliner in a wider sense to accommodate also redundant, symmetry-breaking constraints. This is particularly useful in our context since, for automatically constructed constraints, it is difficult to determine which type the new constraint is, and, from a pragmatic point of view, the distinction does not matter.
2.3 Large Language Models (LLMs)
Large Language Models (LLMs) are advanced AI systems based on the transformer models [?] and trained on vast data sets to produce human-like text and source code in response to instruction prompts [?]. These models, trained on vast amounts of text data, can produce human-like text across various domains and styles [?]. Recent advancements have expanded their capabilities beyond traditional language tasks, including generating, analyzing, and debugging code across multiple programming languages777[????]. In addition to code-related tasks, LLMs have shown promise in mathematical reasoning and problem-solving [??]; they can process and generate mathematical expressions, solve equations, and even assist with proofs. However, it is essential to note that while LLMs can produce seemingly correct mathematical output, their responses should be carefully verified. The models’ performance on these tasks varies, and they may sometimes generate plausible-looking but incorrect solutions [?]. Despite these limitations, the potential applications of LLMs in fields such as computer science, mathematics, and engineering are substantial and continue to expand as the technology evolves.
3 LLM-based Streamliner Generation and Validation
In this section, we discuss the different approaches we devise to generate and test streamliners. The key difference between our approach and the past work by Spracklen et al. [?] is that instead of deriving streamliners by composing and combining atomic constraints, we synthesize entire streamliners in one go. We try several variants of our approach, one which is entirely realtime and requires no precomputation (unlike Spracklen et al. [?]), another which is offline, and one which can directly synthesize combinations of streamliners. We evaluate our approach on benchmark problems from the literature along with modified versions of those problems to give insights into the strengths and weaknesses of our approach.
3.1 Realtime Approach
The realtime approach utilizes the creativity and speed of LLMs to propose and validate streamliners in real time. Such on-the-fly generation is not possible with high-effort streamliner generation that builds a streamliner systematically from elementary steps and takes several CPU days [?].
Our system submits several queries (prompts) to the LLM that result in candidate streamliners and tests the candidate streamliners on a small set of $n_{\text{train}}$ easy training instances that can be solved in less than $t_{\text{train}}$ seconds. Rigorous testing on several problems indicates that a relatively small number of test instances suffices (about 15) and even small and easy test instances (solvable in under 10 seconds with the unstreamlined model) provide a good indication of how well streamliners will work on large and hard instances; we lay out these experiments in more detail in Sections 4.5 and 4.6. We test various strategies for this type of streamliner generation, which we divide into static and adaptive categories. The general strategy can be found in Algorithm 1.
Our realtime approach targets scenarios where a user needs to solve hard constraint problems that typically require hour-long solve times. Rather than directly attempting these hard instances, the user provides a few small instances to our system. This proposes efficient streamliners after interacting with an LLM and a constraint solver within minutes. The user can then run these streamlined versions of the original model in parallel with the unstreamlined version, potentially achieving significant speedup that outweighs the initial streamliner generation time.
Algorithm 1 General Strategy
0:  an unstreamlined MiniZinc model $\mathcal{M}$, a set of training instances $\mathcal{I}$, a prompt $\mathcal{P}$, an LLM $\mathcal{L}$, integers $n,k$
0:  a set of streamliners $\mathcal{T}$
1:  $\mathcal{S}\leftarrow\emptyset$, $\mathcal{S}^{\prime}\leftarrow\emptyset$, $\mathcal{P}^{\prime}\leftarrow\mathcal{P}$
2:  Find the baseline solving times $t_{\mathrm{base}}(I)$ for each instance $I\in\mathcal{I}$, by solving them with the unstreamlined model $\mathcal{M}$.
3:  repeat
4:     Obtain a set $\mathcal{S}^{\prime}$ of $n$ streamliners by prompting the LLM $\mathcal{L}$ with prompt $\mathcal{P}^{\prime}$.
5:     Evaluate the performance of the streamliners from $\mathcal{S}^{\prime}$ in parallel on the training instances using the baseline times $t_{\mathrm{base}}$ as the corresponding timeouts.
6:     Discard from $\mathcal{S}^{\prime}$ the streamliners that produce errors or time out on all the training instances.
7:     $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{S}^{\prime}$
8:     Set $\mathcal{S}^{\prime}\leftarrow\emptyset$ every third iteration.// only relevant for Adaptive Mode
9:     if Adaptive Mode then
10:        Construct the new prompt $\mathcal{P}^{\prime}$ by appending feedback about the streamliners in $\mathcal{S}^{\prime}$ to the prompt $\mathcal{P}$. This feedback includes information such as whether they produced errors or unsatisfiable instances; or if their performance was better or worse than the unstreamlined model.
11:     end if
12:  until 10 minutes have passed.
13:  Find a set $\mathcal{T}$ of $k$ streamliners from $\mathcal{S}$ that collectively achieve the best performance over the training instances.
14:  return  $\mathcal{T}$
3.2 Prompt Engineering
The prompt given in Listing 1, is the prompt we use for most of our experiments. In the beginning, the task is shortly summarized. Then, following the Chain of Thought technique [?], the task is split into single steps that are explicitly explained. In the end, there are some compliance rules to ensure the quality and the correct format of the response. Especially the correct JSON output format is important for the automated tests on the training instances. We decided on this prompt because some preliminary experiments showed good results.
Listing 1 Base Prompt
Objective
Analyze the given MiniZinc code and suggest five additional constraints to enhance the problem-solving process. These constraints can include streamlining, implied, symmetry-breaking, or dominance-breaking constraints.
Steps
1. Analyze Content: Read the provided MiniZinc code. Understand the problem being addressed, including its variables, constraints, and optimization goals. 2. Generate additional Constraints: Based on your analysis, create five unique constraints. These should offer targeted modifications or restrictions designed to reduce the search space effectively. 3. Always return your constraints as a JSON object, adhering to the structure: {“streamliner_1”: “constraint <MiniZinc constraint>”, …, “streamliner_5”: “constraint <MiniZinc constraint>”}. Your final output should exclusively be the JSON object containing the five constraints. 4. As a response, you will get feedback for each constraint. Some constraints might lead to errors, timeouts, or unsatisfiable instances. 5. Use the information provided in the previous step to generate five new, and hopefully better constraints. 6. Repeat steps 3 to 5 multiple times.
Compliance Rules
1. Response Format: Your final output should exclusively be the JSON object containing the five constraints, adhering to the structure: {“streamliner_1”: “constraint <MiniZinc constraint>”, …, “streamliner_5”: “constraint <MiniZinc constraint>”}. Do not forget the semicolon at the end of the constraint! 2. Code Quality: All MiniZinc code provided for the constraints must be syntactically correct and functional. For some functions you may need to include an additional library. 3. Creativity: You’re encouraged to be innovative in proposing constraints, keeping in mind their purpose: to narrow down the search space efficiently without oversimplifying the problem.
3.3 Obfuscation and Disguise
The capabilities of LLMs to generate original ideas versus relying on memorization remains a key debate in AI research [??]. This question is particularly relevant for streamliner generation, as the practical value would be limited if LLMs could only identify streamliners for well-documented problems where such techniques are already published and likely included in their training data. To investigate this capability empirically, we examine two problem variants: disguised and obfuscated versions.
[…]
“The zookeepers at a large zoo are organizing seasonal habitat rotations to encourage socialization among animals. Each season, animals are grouped into mixed-species habitats of a fixed size. The zookeepers want to ensure that no two animals share a habitat more than once during the rotation schedule, while maximizing the number of season the rotations can last. This problem seeks for an allocation of animals to habitats which ensures maximum variety in animal pairings across seasons while maintaining the integrity of habitat sizes.”
Obfuscating a problem is similar to disguising, except that instead of mapping to names with the semantics of a different problem, we map to arbitrary names like id8, id12, and id5, and we strip all comments from the model to avoid revealing the original problem. Thus, for each original problem in our benchmark, we introduce two new problems to our benchmark by disguising and obfuscating the model (.mzn file) and instances (.dzn files) of the original problem. We then generate and evaluate streamliners for these problems using the same prompts as the original problem.
3.4 Combinations of Constraints
Previous work by Spracklen et al. [?] demonstrates that combining multiple streamlining constraints can be more effective than using them individually. To evaluate how combinations of streamliners perform in our StreamLLM approach, we provide another prompt alternative to the base prompt. It was created by modifying the LLM base prompt to output combinations of constraints rather than single constraints, with corresponding adjustments to the required JSON format. We evaluate the results with this alternative prompt in Section 4.9. To establish a baseline for our approach, we evaluate Spracklen et al.’s streamliner combinations on our benchmark instances, with detailed results presented in Subsection 4.10.
3.5 Offline Approach
The offline approach investigates whether extended training time improves streamliner quality. We increase the timeout from 10 minutes to 4 hours and expand the training set from 15 instances solvable within 10 seconds to 35 instances with solving times between 1 and 30 seconds. The process follows the static approach outlined in Algorithm 1. In each iteration, we randomly pick GPT-4o or Claude 3.5 Sonnet, using either the original or combination-generating prompt described in Section 3.4 to leverage the strengths of different LLMs and prompting strategies.
4 Experiments
All instances, MiniZinc models, and the Python implementation of StreamLLM (which includes the prompts) are available on Zenodo [^8] https://zenodo.org/doi/10.5281/zenodo.13331048.
4.1 Setup and Hardware
We use MiniZinc version 2.8.3 and the Chuffed 0.13.1 solver. As LLMs we use GPT-4o (gpt-4o-2024-11-20) and Claude 3.5 Sonnet (claude-3-5-sonnet-20241022) and access them via the openai 1.29.0 and anthropic 0.25.8 packages in Python 3.11.5. We evaluate the running times of the test instances on compute nodes with 2.40GHz, 10-core 2× Intel Xeon E5-2640 v4.
4.2 Benchmark Problems
Although our method can be applied to optimization problems, similar to previous work [?], we focus here on decision problems. This way we can evaluate streamliners only on the basis of their running times. For optimization problems, one needs to consider solution quality as well. We construct a benchmark set of seven constraint satisfaction problems to test the LLM-generated streamliners. This includes four problems which are already considered in previous work by [?], namely Balanced Incomplete Block Design (BIBD) [^2] Problem description and references available at https://csplib.org/Problems [^3] Model from https://csplib.org/Problems, Car Sequencing (CS)22footnotemark: 2444Model from https://github.com/MiniZinc/minizinc-benchmarks, Social Golfers (SG)22footnotemark: 233footnotemark: 3, and Vessel Loading (VL)22footnotemark: 233footnotemark: 3. In addition to that, we include two other well-known problems, Black Hole (BH)22footnotemark: 244footnotemark: 4and Carpet Cutting (CC) [^1] Problem described by Schutt et al. [?]44footnotemark: 4. The main criterion for selecting these six problems is that the MiniZinc model must be readily available. Further filtering is done based on the ease of generation of instances of the desired difficulty. Note that, if present, we strip any redundant constraints, symmetry-breaking constraints, and excessive documentation from the MiniZinc models of these problems. This is the case for CC and SG. The only exception is BIBD, for which, without symmetry-breaking, we are unable to find problem instances of the desired difficulty.
Finally, we added the well-known NP-hard Hypergraph Coloring (HC) problem [??] (described below), for which, to the best of our knowledge, no constraint model is available. The main reason for including this problem is to see whether our approach relies on memorization (e.g., Lee et al. [?] and McCoy et al. [?]).
Hypergraph Coloring (HC): Given integers $c$ and $m$, a set $V$ of vertices, and a set $E$ of hyperedges, where each hyperedge is a subset of $V$, find a coloring of vertices using at most $c$ colors such that no hyperedge is monochromatic, i.e., is only incident to vertices of the same color. Further, the imbalance, i.e., the difference in sizes of the largest and the smallest color class, must not exceed $m$.
4.3 Instance Generation
Figure 1: Histogram showing the number of instances for each problem, sorted by their original running times and partitioned into 20-minute intervals. Further, the total number of instances for each problem is shown next to the problem name in the legend.
Since we work with a broad range of problems, working with instances of similar/uniform difficulty ensures the generality of our findings and prevents bias from creeping in due to any one particular problem. Hence we construct our own data set by generating instances of desired difficulties. For each problem, we create 15 training instances that take less than 10 seconds. Further, we create at least 50 test instances that take between 10 minutes and 2 hours to solve. The 10-minute lower limit is because we want our approach to be comparable in a realtime setting. The 2-hour upper limit is due to limited computational resources.
We generated satisfiable instances of the desired difficulties for all the above problems. Since SG, BH, and BIBD all have relatively simple input specifications, i.e., a small number of integer parameters, we generated instances exhaustively and tested their difficulty. For CC, CS, HC, VL we generated instances using the AutoIG pipeline developed by Dang et al. [?]. This involved coming up with a parameterized Essence specification describing the input space of each problem. We then fix a difficulty window in the form of lower and upper bounds for the MiniZinc running time and then use the automatic algorithm configuration tool irace to find an optimal configuration of those parameters such that the constraint-solving tool Conjure can find an instance that is very likely to fall in the desired difficulty window.
A detailed visualization of the distribution of the original running times (or in other words, difficulties) of the generated instances is shown in Figure 1. The total number of instances for each problem is stated in the legend.
4.4 $k$-Best Selection
As already explained in Section 2.2, there is no guarantee that every satisfiable instance in the original model is also satisfiable in a streamlined model. It also might happen that one streamliner works well on the training instances but is impractical for the large test instances. In order to get more robustness, we decided to not only rely on one streamliner but also to return $k$ streamliners that work best on the training instances. In the experiments, we run the original model and $k=3$ streamlined models in parallel and stop as soon as any of the $k+1$ models has found a solution.
4.5 Preliminary Experiment a
The objective of this experiment is to determine a suitable maximal running time $t_{\text{train}}$ of the unstreamlined model for the training instances. We want to keep $t_{\text{train}}$ as low as possible so that the evaluation can be done quickly but still provide good predictions for the streamliner performance on the significantly larger test instances.
We run the experiment with the problems BIBD, BH, and SG. For each problem, we let the LLM suggest ten streamliners. Then, we pick those three streamliners that perform best on the training instances, and among the three, we pick the one that performs best on the test instances. We use five sets of training instances with an upper bound of $t_{\text{train}}\in\{10,20,30,60\}$ seconds, respectively. The experiments show that $t_{\text{train}}$ has no influence on the one streamliner that was picked in the end, hence we settled on the upper bound $t_{\text{train}}=10$ for the following experiments. To reduce the influence of I/O operations, we require training instances to take at least 1 second to be solved.
4.6 Preliminary Experiment b
Figure 2: Normalized saved times for different number of training instances for three problems
The goal of this experiment is to decide on the number $n_{\text{train}}$ of training instances. More training instances promise better results on the test instances but make the evaluation process longer, which we want to avoid for realtime streamliner generation. Hence we aim at a fair compromise.
We conduct the experiment with the same three problems as in Experiment 1a. For each problem, we generate ten streamliners. For each $n_{\text{train}}\in\{3,5,7,10,20,50\}$, we randomly pick $n_{\text{train}}$ training instances. We determine the combination of three out of the ten streamliners that perform best on the training instances and compute the time this combination saves on the test instances, normalized by the time saved by the virtually best combination of three streamliners. We repeat this 100 times. The box plots in Figure 2 shows the normalized saved running times for different $n_{\text{train}}$. As expected, the larger $n_{\text{train}}$, the more time is saved; setting $n_{\text{train}}=15$ seems a fair compromise for the forthcoming experiments.
4.7 Base Realtime Approach
This is the main experiment where we evaluate the realtime approach. We run StreamLLM on all seven problems, comparing the static and adaptive approach and the two considered LLMs. We conducted the experiment twice to fathom the influence of randomness on the results. Figure 3 reports the average time reduction of both runs.
Figure 3: Percentage of reduction in solving time with both static and adaptive approaches using LLMs Claude and GPT. Each bar and black line denotes, respectively, the mean and standard deviation of two runs.
Overall, the realtime approach achieves very encouraging results. Some runs archive reduction in running times of 98% and more. This is the case for the CS, the SG, and the HC problem. In some cases, the streamlined models could even solve most of the test instances in less than a second. The good performance on the HC problem suggests that the LLMs do not just copy and paste known streamliners from literature, but are also capable of dealing with new problems. For a few of the problems, most noticeably the BIBD problem, the reduction is only moderate. This is not so surprising considering that the model for BIBD already includes symmetry-breaking constraints and so it turns out to be more challenging to find significant time savings. It is worth noting that BIBD also sees very minor improvement in the approach by Spracklen et al. [?].
Overall, Claude works slightly better than GPT. Comparing the adaptive and static approaches, there is no clear winner. This might be because adaptive limits the creativity of the LLMs.
Figure 4: Solving time reduction with respect to the original solving time when including streamliner generation time in the total streamlined running time.
Our analysis thus far compared the time reduction of streamlined models over the unstreamlined model, but did not consider the time StreamLLM spent on streamliner generation. To revisit the realtime scenario as sketched in the introduction, Figure 4 shows the percentage of saved time when including the streamliner generation time in the running time of the streamlined model [^9] Since there are only a few instances where the unstreamlined model takes between 100 and 120 minutes, we removed this interval.. As expected, the saved time for instances that take less than 20 minutes is relatively poor since more than half of the time is spent on streamliner generation. For larger instances, however, the generation time is relatively insignificant, and StreamLLM shows a remarkably strong performance on most of the problems, with the exception of the BIBD problem. Note that, the BH problem has no instances that take more than 60 minutes.
4.8 Obfuscation and Disguise
Figure 5: Percentage of reduction in solving time for all seven original problems, as well as their disguised and obfuscated versions, with both static and adaptive approaches using LLMs Claude and GPT. The bars and black lines denote the means and standard deviations, respectively.
This experiment evaluates how obfuscation and disguise affect the performance of our StreamLLM approach. Figure 5 shows the results. It is interesting to see is that for GPT, our approach performs, on average, slightly better on disguised problems than on the original problems. For Claude, it performs slightly better on the original problems. In any case, the good results on the disguised problems suggest that the LLMs do not only copy and paste online available streamliners for known problems but can make sense of the underlying problem and can find streamliners for new problems. For the obfuscated problems, our approach performs, on average, worse than on the original and disguised problems. This suggests that LLMs prefer semantically richer language over pure abstract terminology to make better sense of variables and constraints.
4.9 Combinations
We evaluate streamliner combinations on the problems BH, CC, and VL, which we selected because they exhibited moderate performance improvements with individual streamliners, neither the dramatic gains seen in problems like SG nor the minimal improvements observed with BIBD. As shown in Figure 6, combining streamliners proves more effective than individual constraints for BH and VL while slightly reducing performance for CC.
Figure 6: Percentage of reduction in solving time obtained by combinations of streamliners with both static and adaptive approaches using LLMs Claude and GPT. The bars and black lines denote the means and standard deviations, respectively.
4.10 Comparison with Previous Work
We replicate a subset of Spracklen et al.’s [?] experiments on our benchmark instances to establish a baseline for comparing our base approach against. However, this comparison should be interpreted cautiously due to the following significant differences between the approaches: (i) the technical pipelines differ—their method uses Conjure and Essence for solving and modeling, whereas we employ MiniZinc for both purposes; (ii) the benchmark problems used in their experiments differ from our benchmark set.
We used the top-performing streamliners from the repository published by Spracklen et al. [^10] https://github.com/stacs-cp/automated-streamliner-portfolios as the initial candidate pool and evaluated them on our benchmark instances. BIBD, CS, SG, and VL are the only problems that appear in both benchmarks. However, the Conjure model for VL is significantly different from our MiniZinc model, making it incompatible with our input instances. Thus, we only compare the remaining three problems. We first evaluated the initial candidates on our training instances to determine the three best streamliners for each problem, which we then evaluated on the test instances. The best triples saved 3%, 17%, and 21% time for the BIBD, CS, and SG problems, respectively. Only three of the nine streamliners from these triples could satisfy all the test instances.
We see two reasons for these savings being much lower than those reported by Spracklen et al.: (i) we consider all instances towards the total savings, while they only average over the improved instances; (ii) the difficulty of our benchmark instances was calibrated for MiniZinc, which need not necessarily align with the difficulty for Closure. Finally, it is also worth noting, for pragmatic reasons, we compared their approach against our base approach, which operates in real time and does not generate and train streamliners offline like their method. Furthermore, their approach allows streamliners to be combinations of constraints, while our base approach considers only individual constraints.
4.11 Offline Approach
We conduct the offline approach on the BH and CC problems, which show a middle-ground performance in the main experiment, running each problem twice to assess the impact of randomness. Figure 7 presents the performance of the current best triple during training and its corresponding performance on test instances. The strong correlation between training and test instance trends validates our approach of inferring streamliner quality from smaller training instances. The performance graph shows significant quality improvements in the early stages before gradually leveling off. By the end of training, all runs achieve triples of streamliners that save over 90% of running time on test instances, demonstrating that this offline scenario with extended training times yields better results than the realtime approach. Among the top triples from each run, combinations make up slightly more than half of the streamliners, with Claude generating two-thirds and GPT-4 contributing one-third. This distribution suggests that the diversity achieved through multiple prompts and LLMs enhances overall results.
Figure 7: Percentage of reduction in solving time on both training and test instances with respect to the length of the training phase for two runs each on two problems.
5 Analysis of the Streamliners
In this section, we analyze the relationship between the individual streamliners within the best-performing triple for each problem. These triples are selected based on their performance in the main experiment outlined in Section 4.7.
Table 1 shows the contribution of each of the streamliners and the original model, i.e., the percentage of test instances it performs best. The table also shows the percentage of unsatisfiable test instances for each streamliner. Further, one can see whether a streamliner is an implied constraint, a symmetry-breaking constraint, or does not preserve satisfiability.
It is interesting to see that for the different problems, the contribution of streamliners is distributed very differently. For the HC problem, one of the streamliners contributes to 98% of the instances, while the other two have almost no impact. In contrast to this, for the VL problem, the contribution of the three streamliners, as well as the original model, are distributed almost evenly. Further, one can see that for some of the problems, the original model has no contribution at all, which means that for each instance there is at least one of the streamliners in the triple that performs better than the original model.
It is worth noting that for some of the triples, the streamlined models make none or only a few of the instances unsatisfiable, whilst others have rather high percentages of unsatisfiable instances. Especially for the BH and the BIBD problem, even the streamliner with the highest contribution makes about one-third of the instances unsatisfiable. Of the 21 total streamliners mentioned in Table 1, 10 streamliners managed to satisfy all the test instances. We further manually inspected these 10 streamliners to find that 7 of them are indeed solution preserving, i.e., either implied or symmetry-breaking. All the 21 streamliners are listed in the appendix.
Problem Contribution (%) Unsat (%) Type
orig. s1 s2 s3 s1 s2 s3 s1 s2 s3
BH 11 60 26 2 32 13 77 - - -
CC 0 85 12 4 2 2 44 - - -
BIBD 7 56 31 6 36 0 79 - u -
SG 0 74 25 2 0 5 0 s* - s*
VL 25 29 24 22 0 0 0 i i i
CS 0 84 12 3 0 5 0 i - u
HC 0 98 1 1 0 0 1 i u -
Table 1: Percentages of contribution and percentage of unsatisfiable instances for all streamliners of the best-performing triples for all problems. The type column indicates whether the streamliner is an implied constraint (i), a symmetry-breaking constraint (s), or a constraint that does not preserve satisfiability (-). A (*) indicates that the type has not been formally verified.
6 Conclusion and Future Work
Our StreamLLM system and its analysis demonstrate that LLMs can effectively generate streamlining constraints for constraint satisfaction problems. Our experiments across seven diverse problems show impressive runtime reductions compared to unstreamlined models. The system can generate effective streamliners within minutes using only small training instances, making it practical even for problems with hour-long solving times. StreamLLM performs significantly better on the original and the disguised problem variants than on the obfuscated ones. This suggests a similarity to human experts, who naturally leverage the semantic context.
Our work gives rise to several promising research directions. (i) Knowledge distillation could compress the LLM components into smaller, locally deployable language models specifically tuned for streamliner generation. (ii) The observation that disguised problems sometimes yield better results suggests systematically exploring problem translations as a meta-optimization strategy. (iii) Developing LLM-supported formal verification methods for streamliner soundness could apply StreamLLM to unsatisfiable instances. And finally, (iv) combining LLM-generated streamliners with evolutionary optimization techniques could improve their effectiveness while maintaining the rapid generation capabilities demonstrated in this work.
Appendix: Streamliners
Black Hole
s1: constraint x[26] == 26;
s2: constraint forall(i in 1..51) (x[i] != x[i+1] + 1);
s3: constraint x[52] == 52;
Carpet Cutting
s1: constraint forall(i in Stairs, j in st_rec_ids[i])(st_rec_x[j] mod st_rec_len[j] = 0);
s2: constraint forall(i in Stairs, j in st_rec_ids[i])(st_rec_y[j] mod st_rec_wid[j] = 0);
s3: constraint forall(i in Rooms, j in rm_rec_ids[i])(rm_rec_x[j] mod MinRmRecSize = 0);
Balanced Incomplete Block Design
s1: constraint forall(i in rows)(sum(j in cols where j mod 2 = 1)(bool2int(m[i, j])) >= r div 2);
s2: constraint forall(i in rows)(sum(j in cols)(bool2int(m[i, j])) >= k div 2);
s3: constraint forall(i in rows, j in cols where i = j)(m[i, j] = false);
Social Golfers
s1: constraint forall(g in Golfer)(assign[g,2] = ((g-1) + n_per_group) mod n_groups + 1);
s2: constraint forall(w in Week, gr in Group)(assign[gr,w] = gr);
s3: constraint forall(g in Golfer)(assign[g,1] = (g-1) div n_per_group + 1);
Vessel Loading
s1: constraint forall(c in Containers) (Left[c] <= deck_width - width[c]);
s2: constraint max([Right[c] | c in Containers]) <= deck_width;
s3: constraint forall(c1, c2 in Containers where c1 < c2)(Left[c1] <= deck_width - width[c1] \/ Left[c2] <= deck_width - width[c2]);
Car Sequencing
s1: constraint forall(i in options)(count(j in 1..n_cars)(b_seq_confs[i,j] = 1) = sum(c in configurations)(confs[c,i] * n_cars_by_confs[c]));
s2: constraint forall(i in 1..n_cars-1)(seq_confs[i] >= seq_confs[i+1] -> forall(j in options)(b_seq_confs[j,i] >= b_seq_confs[j,i+1]));
s3: constraint forall(i in 1..n_cars-2)(seq_confs[i] = seq_confs[i+1] -> seq_confs[i+1] != seq_confs[i+2]);
Hypergraph Coloring
s1: constraint sum(c in Color)(num_vertices_of_color[c]) = num_vertices;
s2: constraint forall(c in Color)(num_vertices_of_color[c] <= (num_vertices div num_colors) + (max_imbalance div 2));
s3: constraint forall(v1, v2 in Vertex where v1 < v2)(if exists(e in Hyperedge)(incidence[e,v1] = 1 / incidence[e,v2] = 1) then coloring[v1] <= coloring[v2] endif);
Acknowledgments
The authors thank Carlos Ansótegui for helpful discussions. This research is supported by the Austrian Science Funds (FWF), projects 10.55776/COE12, 10.55776/P36688, and 10.55776/P36420. Part of this work was carried out while the third author was a visiting researcher at the Simons Institute for the Theory of Computing, Berkeley.